Please refresh for the latest information on this outage.   Latest updates will be posted at the top.

5.34pm
The servers are back up.  The issue was a switch issue.  The technician in the data centre is flagging this for the data centre networking team in the morning as what happened should not have happened.  We were told the switch port lost its VLAN configuration meaning none of our IP ranges were working on the server.  We're sorry for this outage and we hope our announcements have been helpful.  If you have any more questions please reach out to our team via a support ticket.

5.20pm
We are in the server but there is no networking access.  We cannot access the gateway of any of our IP ranges.  We've tried adding different IPs but none are pinging.  This very much sounds like a data centre networking issue and we're putting this issue back over to them now with high priority.

5.10pm

Our remote hands are now at the server and have told us the server is at the login prompt but there is no external networking.  They are now connecting the KVM remote stand so we can log into the server and determine why external networking is broken.  We're sorry about this and we will have this resolved as quickly as possible.

4.25pm

Our alerting system reported that the node was down.  We checked and confirmed this and issued a reboot request. As the reboot did not bring the server back on line we have asked the remote hands in the data centre to connect a remote stand (KVM) to the server.  They estimate this will be available by 5.10pm and once we have that our team will be able to log in and check what is going on.

Thursday, November 23, 2017

« Back